Ali deposites $1750.00 in a registered account on march 1st 1988 at 10% compounded quarterly.Subsequently the interest rate was changed to 12% compounded monthly on september 1st,1990 and to 11% compunded semi-annually on june 1st,1992.What will the value of the deposit be on december 1st,1998,if no further changes in interest rates are changed?
Answer by rfer(16322)

A=P(1+r/n)^rn
3-1-88 $1750 at 10% comp quarterly
9-1-90 $2106 at 12% comp monthly
6-1-92 $2569.72 at 11% comp semi annually
12-1-98 total = $3738.12
------------------------------------
A=1750(1+.1/4)^7.5=1750(1.025)^7.5=1750(1.2035)=2106
A=2106(1+.12/12)^20=2106(1.01)^20=2106(1.220)=2569.72
A=2569.72(1+.11/2)^7=2569.72(1.055)^7=2569.72(1.4547)=3738.12